Sean has one less than twice the number of stickers that Jason has. Rick has 53 more stickers than Jason has. If Sean gives Jason 25 stickers, Sean and Jason will have the same number of stickers. What is the total number of stickers the three boys have? With exact solution.
Answer by JulietG(1812)

Here is the information you're given:
S = 2J-1
R = J+53
S-25 = J+25, or J=S-50
.
Take the value of J from the bottom equation and plug it into the top equation.
S = 2(S-50)-1
S = 2S-100-1
S = 2S-101
Add 101 to each side
S + 101 = 2S
Subtract S from each side
101 = S
.
If Sean has 101 stickers, then Jason has 51 stickers. If Jason has 51 stickers, then Rick has 104 stickers.
.
S + J + R = x
101 + 51 + 104 = 256